# Resource-Efficient Hybrid Quantum-Classical Simulation Algorithm

### v1

@@ Acknowledgements.
Appendix A Modified Hadamard Test
The standard Hadamard Test estimates the real and imaginary components of $\bra{\psi}\hat{U}\ket{\psi}$ using an ancilla qubit and an $N$-system-qubit unitary $U$ gate controlled by the ancilla qubit. This creates a problem: parallel gates in the gate decomposition of $U$ must be serialized when gate control by the ancilla qubit is applied. We shall show that this problem can be solved using more ancilla qubits. As a simple example, let $\hat{U}{=}\hat{U}_{1}{\otimes}\hat{U}_{2}$ where $\hat{U}_{1}$ and $\hat{U}_{2}$ act on two separate sets of system qubits. Then, we may consider the following modified Hadamard Test that uses two ancilla qubits, as shown in Fig. 4.
@@ Appendix A Modified Hadamard Test
Appendix B Quantum Resource-Efficient Regime
@@
We assume the system Hamiltonian $\hat{H}$ and observable $\hat{O}$ both decompose into a linear combination of $L$ Pauli strings $\hat{P}_{l}{=}\otimes^{N}_{j=1}\hat{\sigma}_{j}$, where $\hat{\sigma}_{j}{\in}\{\hat{I}_{j},\hat{X}_{j},\hat{Y}_{j},\hat{Z}_{j}\}$. The quantum runtime complexity of existing quantum algorithms that simulates the $e^{-iH\Delta t}$ time propagator up to a time evolution error $\epsilon{\geq}\left\|e^{i\hat{H}\Delta t}{-}\hat{U}\right\|$ are upper bounded by $\mathcal{O}(\textrm{Poly}_{\hat{U}}(L,\epsilon^{-1},\Delta t))$ as shown in Table 1. Thus, performing the time evolution for a simulation time of $j\Delta t$, where $j{\in}\mathcal{Z}$ is a non-negative integer, incurs a quantum runtime complexity of $j\cdot\mathcal{O}(\textrm{Poly}_{\hat{U}}(L,\epsilon^{-1},\Delta t))$ from applying the $e^{-iH\Delta t}$ time propagator $j$ times on the initial state.
Algorithms Simulating $e^{-iHt}$ Quantum runtime complexity
@@
Qubitization [19] $\mathcal{O}\left[t\|\hat{H}\|_{\max}+\log(1/\epsilon)/\log\log(1/\epsilon)\right]$
Linear Combination of Unitaries [15] $\mathcal{O}\left[t\lambda+\log(t\lambda/\epsilon)/\log\log(t\lambda/\epsilon)\right]$
@@
Table 1: The quantum runtime or gate complexity of various quantum algorithms simulating $e^{-iHt}$ time propagator, that are upper bounded by $\mathcal{O}(\textrm{Poly}_{\hat{U}}(L,\epsilon^{-1},t))$, for a time-independent Hamiltonian $\hat{H}$ and up to a simulation time $t$ [6]. Here $\|\hat{H}\|_{\max}$ is the largest absolute element of $\hat{H}$ and $\lambda{=}\sum_{l}p_{l}$ is the sum of all Pauli coefficients of the Pauli-form of Hamiltonian $\hat{H}$.
Due to wavefunction collapse and no-fast-forwarding theorem, the standard method involves preparing $L$ copies of time-evolved states at times $\{j\Delta t|j{=}0,1,\ldots,\frac{T}{\Delta t}\}$ for Pauli measurements. Since, in the main text, we assume access to a $2N$-qubit quantum computer, we may parallelize the time-evolution to reduce the overall runtime by a factor of 2. Therefore, the overall quantum runtime complexity of the standard method is
@@
For the QAS, the quantum computer is used only for estimating the real and imaginary components of the overlap $\boldsymbol{F}$, Hamiltonian $\boldsymbol{H}$, and observable $\boldsymbol{O}$ matrices, as shown in the main text section II.1. We choose $n$ time-evolved states |$\ket{\psi_{j}}=e^{-i\hat{H}s_{j}}\ket{\psi_{0}}$ as a basis, where $j{=}\{0,1\ldots,n{-}1\}$ and parameter times $0{=}s_{0}{<}s_{1}{<}{\ldots}{<}s_{n-1}{\leq}T$ are not longer than the total simulation time. As a result, the aforementioned matrix elements consist of the quantities $F_{jk}{=}\braket{\psi_{0}}{e^{i\hat{H}\Delta s_{jk}}}{\psi_{0}}$ and $P_{jkl}{=}\braket{\psi_{0}}{\hat{P}_{l}e^{i\hat{H}\Delta s_{jk}}}{\psi_{0}}$ where $k{=}\{0,1,\ldots,n{-}1\}$, $\Delta s_{jk}{=}s_{j}{-}s_{k}$ are the parameter time differences and $l{=}\{0,1,\ldots,2L\}$. We consider using the modified Hadamard test from Appendix A to estimate them. It requires a controlled-$e^{-iH\Delta t}$ time propagator that has a longer quantum runtime complexity than a standard $e^{-iH\Delta t}$ time propagator by at most a constant factor of $\gamma$. Thus, the overall quantum runtime complexity of the QAS is
@@
Appendix C Sampling Basis State Overlaps, Hamiltonian and Observable Elements.
@@
Appendix D Energy Dynamics of Helium Atom and Hydrogen Molecule
@@
Appendix E Variance of Observed Quantities against Number of Shots
@@
E.1 Helium Atom Case
@@
E.2 Hydrogen Molecule Case
@@
Appendix F First Order Trotterized Time-Evolved Basis States
@@
Appendix G Linear Dependence of the Time-Evolved Basis

### v2

@@ Acknowledgements.
Appendix A Derivation of Parameter Evolution
We present the derivation of the parameter evolution, adapted from Bharti and Haug [40], and Yuan et al. [44]. In the main text, we define the time-evolution state ansatz as
|$\ket{\psi(t)}=\sum_{j=0}^{n_{\textrm{basis}}-1}\alpha_{j}(t)\ket{\psi_{j}}.$ (13)
The McLachlan’s variational principle [53] states that the quantum time-evolution always minimizes the square root error overlap between the left and right hand side states of the Schrödinger equation,
$\delta_{\boldsymbol{\alpha}}\|(i\hbar\partial_{t}-\hat{H})\ket{\psi(t)}\|=0.$ (14)
where error overlap $\epsilon$,
$\epsilon$ $=$ $\|(i\hbar\partial_{t}-\hat{H})\ket{\psi(t)}\|^{2}$ (15)
$=$ $[(i\hbar\partial_{t}-\hat{H})\ket{\psi(t)}]^{\dagger}[(i\hbar\partial_{t}-\hat{H})\ket{\psi(t)}]$ (17)
$=$ $\left\{\sum_{j,k=0}^{n_{\textrm{basis}}-1}[\partial_{\alpha^{*}_{j}}\bra{\psi(t)}][\partial_{\alpha_{k}}\ket{\psi(t)}]\dot{\alpha}_{j}^{*}\dot{\alpha}_{k}\right\}+i\left\{\sum_{j=0}^{n_{\textrm{basis}}-1}[\partial_{\alpha^{*}_{j}}\bra{\psi(t)}]\hat{H}\ket{\psi(t)}\dot{\alpha}_{j}^{*}\right\}$
$\quad-i\left\{\sum_{k=0}^{n_{\textrm{basis}}-1}\bra{\psi(t)}\hat{H}[\partial_{\alpha_{k}}\ket{\psi(t)}]\dot{\alpha}_{k}\right\}+\bra{\psi(t)}\hat{H}^{2}\ket{\psi(t)}.$
Since,
$\partial_{\alpha_{j}}\ket{\psi(t)}=\ket{\psi_{j}},$ (18)
the error overlap $\epsilon$ in Eq. (17) simplifies to
$\epsilon$ $=$ $\left\{\sum_{j,k=0}^{n_{\textrm{basis}}-1}\braket{\psi_{j}}{\psi_{k}}\dot{\alpha}_{j}^{*}\dot{\alpha}_{k}\right\}+i\left\{\sum_{j=0}^{n_{\textrm{basis}}-1}\bra{\psi_{j}}\hat{H}\ket{\psi(t)}\dot{\alpha}_{j}^{*}\right\}$ (19)
$\quad-i\left\{\sum_{k=0}^{n_{\textrm{basis}}-1}\bra{\psi(t)}\hat{H}\ket{\psi_{k}}\dot{\alpha}_{k}\right\}+\bra{\psi(t)}\hat{H}^{2}\ket{\psi(t)}$
$=$ $\dot{\boldsymbol{\alpha}}^{\dagger}\boldsymbol{F}\dot{\boldsymbol{\alpha}}-i\dot{\boldsymbol{\alpha}}^{\dagger}\boldsymbol{H}\boldsymbol{\alpha}+i\boldsymbol{\alpha}^{\dagger}\boldsymbol{H}\dot{\boldsymbol{\alpha}}+\boldsymbol{\alpha}^{\dagger}(\boldsymbol{H^{2}})\boldsymbol{\alpha},$ (20)
where have used the following matrices
$F_{jk}$ $=$ $\braket{\psi_{j}}{\psi_{k}},$ (21)
$H_{jk}$ $=$ $\bra{\psi_{j}}\hat{H}\ket{\psi_{k}},$ (22)
$(\boldsymbol{H^{2}})_{jk}$ $=$ $\bra{\psi_{j}}\hat{H}^{2}\ket{\psi_{k}}.$ (23)
As the variation of the square root error overlap is equivalent to the variation of the error overlap up to a constant factor, we may focus on variation of the error overlap instead, then we have
$\delta_{\alpha_{k}}\epsilon$ $=$ $\left(\sum_{j=0}^{n_{\textrm{basis}}-1}[\partial_{\alpha^{*}_{j}}\bra{\psi(t)}][\partial_{\alpha_{k}}\ket{\psi(t)}]\dot{\alpha}_{j}+i[\partial_{\alpha_{k}}\ket{\psi(t)}]\hat{H}\ket{\psi(t)}\right)\delta\dot{\alpha}^{*}_{k}$ (24)
$\quad+\left(\sum_{j=0}^{n_{\textrm{basis}}-1}[\partial_{\alpha^{*}_{j}}\bra{\psi(t)}][\partial_{\alpha_{k}}\ket{\psi(t)}]\dot{\alpha}^{*}_{j}-i[\partial_{\alpha_{k}}\ket{\psi(t)}]\hat{H}\ket{\psi(t)}\right)\delta\dot{\alpha}_{k}$
$=$ $(\boldsymbol{F}\dot{\boldsymbol{\alpha}}+i\boldsymbol{H}\boldsymbol{\alpha})\delta\dot{\alpha}^{*}_{k}+(\boldsymbol{F}\dot{\boldsymbol{\alpha}}^{*}-i\boldsymbol{H}\boldsymbol{\alpha})\delta\dot{\alpha}_{k}.$ (25)
Thus, the parameter evolution is
$\boldsymbol{F}\dot{\boldsymbol{\alpha}}=-i\boldsymbol{H}\boldsymbol{\alpha},$ (26)
which minimizes the error overlap $\epsilon$ in Eq.(20) throughout the entire evolution,
$\epsilon_{\textrm{min}}=\boldsymbol{\alpha}^{\dagger}(\boldsymbol{H^{2}})\boldsymbol{\alpha}-\dot{\boldsymbol{\alpha}}^{\dagger}\boldsymbol{F}\dot{\boldsymbol{\alpha}}.$ (27)
Appendix B Modified Hadamard Test
The standard Hadamard Test estimates the real and imaginary components of $\bra{\psi}\hat{U}\ket{\psi}$ using an ancilla qubit and an $N$-system-qubit unitary $\hat{U}$ gate controlled by the ancilla qubit. This creates a problem: parallel gates in the gate decomposition of $\hat{U}$ must be serialized when gate control by the ancilla qubit is applied. We shall show that this problem can be solved using more ancilla qubits. As a simple example, let $\hat{U}{=}\hat{U}_{1}{\otimes}\hat{U}_{2}$ where $\hat{U}_{1}$ and $\hat{U}_{2}$ act on two separate sets of system qubits. Then, we may consider the following modified Hadamard Test that uses two ancilla qubits, as shown in Fig. 4.
@@ Appendix B Modified Hadamard Test
Appendix C Quantum Resource-Efficient Regime
@@
We assume the system Hamiltonian $\hat{H}$ and observable $\hat{O}$ both decompose into a linear combination of $L$ Pauli strings $\hat{P}_{l}{=}\otimes^{N}_{j=1}\hat{\sigma}_{j}$, where $\hat{\sigma}_{j}{\in}\{\hat{I}_{j},\hat{X}_{j},\hat{Y}_{j},\hat{Z}_{j}\}$, which may or may not share the same Pauli string. The quantum runtime complexity of existing quantum algorithms that simulates the $e^{-i\hat{H}\Delta t}$ time propagator up to a time evolution error $\epsilon{\geq}\left\|e^{i\hat{H}\Delta t}{-}\hat{U}\right\|$ are upper bounded by $\mathcal{O}(\textrm{Poly}_{\hat{U}}(L,\epsilon^{-1},\Delta t))$ as shown in Table 1. Thus, performing the time evolution for a simulation time of $j\Delta t$, where $j{\in}\mathcal{Z}$ is a non-negative integer, incurs a quantum runtime complexity of $j\cdot\mathcal{O}(\textrm{Poly}_{\hat{U}}(L,\epsilon^{-1},\Delta t))$ from applying the $e^{-i\hat{H}\Delta t}$ time propagator $j$ times on the initial state.
Algorithms Simulating $e^{-i\hat{H}t}$ Quantum runtime complexity
@@
Qubitization [19] $\mathcal{O}\left[t\lambda+\log(1/\epsilon)/\log\log(1/\epsilon)\right]$
Linear Combination of Unitaries [15] $\mathcal{O}\left[t\lambda\log(t\lambda/\epsilon)/\log\log(t\lambda/\epsilon)\right]$
@@
Table 1: The quantum runtime or gate complexity of various quantum algorithms simulating $e^{-i\hat{H}t}$ time propagator, that are upper bounded by $\mathcal{O}(\textrm{Poly}_{\hat{U}}(L,\epsilon^{-1},t))$, for a time-independent Hamiltonian $\hat{H}$ and up to a simulation time $t$ [6]. Here $\|\hat{H}\|_{\max}$ is the largest absolute element of $\hat{H}$ and $\lambda{=}\sum_{l}p_{l}$ is the sum of all Pauli coefficients of the Pauli-form of Hamiltonian $\hat{H}$.
Due to wavefunction collapse and no-fast-forwarding theorem, the standard method involves preparing $L$ copies of time-evolved states at times $\{j\Delta t|j{=}0,1,\ldots,\frac{T}{\Delta t}\}$ for Pauli measurements. Since, in the main text, we assume that QAS have access to a $2N$-qubit quantum computer, we may parallelize the the standard method by running two independent time-evolution at a time, reducing the overall runtime by a factor of 2. Therefore, the overall quantum runtime complexity of the standard method is
@@
For the QAS, the quantum computer is used only for estimating the real and imaginary components of the overlap $\boldsymbol{F}$, Hamiltonian $\boldsymbol{H}$, and observable $\boldsymbol{O}$ matrices, as shown in the main text section II.1. We choose $n$ time-evolved states |$\ket{\psi_{j}}=e^{-i\hat{H}s_{j}}\ket{\psi_{0}}$ as basis, where $j{=}\{0,1\ldots,n{-}1\}$ and parameter times $0{=}s_{0}{<}s_{1}{<}{\ldots}{<}s_{n-1}{\leq}T$ are not longer than the total simulation time. As a result, the aforementioned matrix elements consist of the quantities $F_{jk}{=}\braket{\psi_{0}}{e^{i\hat{H}\Delta s_{jk}}}{\psi_{0}}$ and $P_{jkl}{=}\braket{\psi_{0}}{\hat{P}_{l}e^{i\hat{H}\Delta s_{jk}}}{\psi_{0}}$ where $k{=}\{0,1,\ldots,n{-}1\}$, $\Delta s_{jk}{=}s_{j}{-}s_{k}$ are the parameter time differences and $l{=}\{0,1,\ldots,2L\}$. In total, there are $n^{2}{+}2Ln^{2}{=}\mathcal{O}(Ln^{2})$ such quantities combined. We consider using the modified Hadamard test from Appendix B to estimate them. It requires a controlled-$e^{-i\hat{H}\Delta t}$ time propagator that has a longer quantum runtime complexity than a standard $e^{-i\hat{H}\Delta t}$ time propagator by at most a constant factor of $\gamma$. Thus, the overall quantum runtime complexity of the QAS is
@@
Appendix D Sampling Basis State Overlaps, Hamiltonian and Observable Elements.
@@
Appendix E Energy Dynamics of Helium Atom and Hydrogen Molecule
@@
Appendix F Variance of Observed Quantities against Number of Shots
@@
F.1 Helium Atom Case
@@
F.2 Hydrogen Molecule Case
@@
Appendix G First Order Trotterized Time-Evolved Basis States
@@
Appendix H Linear Dependence of the Time-Evolved Basis